OpenStudy (anonymous):

Solve the given optimization problem by using substitution. HINT [See Example 1.] Find the minimum value of f(x, y, z) = x^(2) + y^(2) + z^(2) − 7 subject to z = y.

ganeshie8 (ganeshie8):

use the hint ?

OpenStudy (anonymous):

doesn't really help

OpenStudy (anonymous):

its the same thing for every problem

ganeshie8 (ganeshie8):

substitute z=y in the given function, that eliminates z from the function

ganeshie8 (ganeshie8):

f(x, y, z) = x^(2) + y^(2) + z^(2) − 7 put z = y f(x, y) = x^2 + y^2 + y^2 - 7 = x^2 + 2y^2 - 7

OpenStudy (anonymous):

so fx = 2x fy = 4y

ganeshie8 (ganeshie8):

oh u wana use lagrange ? its a overkill for this problem..

ganeshie8 (ganeshie8):

think a bit, whats the minimum value of this ? can the function get any lower than -7 ?

OpenStudy (anonymous):

no, so -7 is the min?

ganeshie8 (ganeshie8):

since x^2+2y^2 is always positive, the minimum value is 0 for this. so, when x=y=z=0, the function has a minimum value of -7

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

ohh okay that makes sense

ganeshie8 (ganeshie8):

good, if u wana carry out lagrange... u may do that as well... guess ur professor wants u show that work

OpenStudy (anonymous):

yeah he taught us through that way but he hasn't done examples that are on the homework thats why I'm confused